# Interpreting GNN-based IDS Detections Using Provenance Graph Structural Features

### v5

@@ 8 Conclusion
We introduced ProvExplainer that aims to bridge a critical gap in explainable ML for cybersecurity by providing transparent, human-interpretable insights into GNN-based decisions for provenance-based intrusion detection. By integrating security-aware subgraph patterns and structural graph features, ProvExplainer enables effective instance-level explanations that are sensitive to both local and global graph properties. This dual-level interpretability is instrumental in addressing the opacity that limits the adoption of GNN-based models in critical security environments.
Through rigorous evaluation on real-world APT and Fileless Malware  datasets, ProvExplainer demonstrates superior fidelity and accuracy in its explanations, significantly outperforming existing SOTA GNN explainers across $\text{fidelity}^{+}$ , $\text{fidelity}^{-}$ , precision, and recall metrics. ProvExplainer increases $\text{fidelity}^{+}$  and decreases $\text{fidelity}^{-}$  by 29% and 12%, and precision and recall by 27% and 25% over SOTA GNN explainers with respect to attack documentation prepared by security vendors and dataset authors. Importantly, while ProvExplainer structure-agnostic design allows for easy extension with additional security-aware features. This makes ProvExplainer not only an advancement in explainable AI for security applications but also a scalable and adaptable framework for future research and deployment in real-world security contexts.
[…]
@@ Appendix A Appendix
A.1 Implementation
Our data collection module can operate on Windows systems by using the Windows ETW [winETW] and on Linux system by using Linux auditd [audit] frameworks to collect relevant system calls regarding files, processes and network sockets. These include system calls for (1) file operations (e.g., read(), write(), unlink()), (3) network socket operations (e.g., connect(), accept()), (4) process operations (e.g., create(), exec(), and exit()). The system-level data is stored in a PostgresSQL database. Table I summarizes the event collection schema  for our in-house deployment.
ProvExplainer is implemented using python and graph generation framework is implemented in java. The GNN training and evaluation pipelines are implemented using Deep Graph Library (DGL) [dgl] framework, and the surrogate DTs are implemented using sklearn [scikit].
A.2 Dataset Statistics
@@
Applications # of Benign Graphs # of Anomaly Graphs Avg # of Benign Nodes / Edges Avg # of Anomaly Nodes / Edges
@@
Fileless Malware from [survivialism2021sp]
@@
Anomaly Detection Dataset. The dataset statistics for the anomaly detection dataset is seen in Table IV. The Fileless Malware samples were downloaded from [virustotal] and selected from recent studies, [grammatikakis2021csr, kuchler2021does, avllazagaj2021malware, survivialism2021sp]. The benign provenance graphs for the anomaly detection dataset were sourced from the DARPA dataset, which contained 5,695 benign graphs with an average of 669.81 vertices and 982.18 edges (Table IV); the Fileless Malware  dataset, which contained 19,422 benign graphs with an average of 60.44 vertices and 54.15 edges; and the APT dataset, which contained 6,291 benign graphs with an average of 55.94 vertices and 83.29 edges. The corresponding anomaly graphs are sourced from the same datasets: The DARPA dataset contained 30 anomalous graphs with an average of 777.08 vertices and 1011.19 edges (Table IV); the Fileless Malware  contained 1,043 anomalous graphs with an average of 63.72 vertices and 89.93 edges; and the APT dataset contained 2,928 graphs with an average of 58.38 vertices and 51.69 edges.
A.3 Ablation Study (using GraphSAGE)
Table V: Agreement of surrogate DTs with the GraphSAGE  model across different feature subsets. The best feature subsets are highlighted.
Dataset Top-$K$ Subgraphs All Top 5 Subgraphs Graph Structural Features All Graph Structural Feat. All Features
1 2 3 4 5 Deg. Cen. Clos. Cen. Bet. Cen Eig. Cen. Clus. Coeff. Clus. Tri.
APT Dataset from [mukherjee2023sec]
Enterprise 0.91 (-0.03) 0.91 (-0.03) 0.91 (-0.03) 0.91 (-0.03) 0.91 (-0.03) 0.91 (-0.03) 0.91 (-0.03) 0.92 (-0.02) 0.92 (-0.02) 0.92 (-0.02) 0.92 (-0.02) 0.94 (+0.00) 0.94 (+0.00) 0.94
Supply-Chain 0.77 (-0.20) 0.88 (-0.09) 0.89 (-0.08) 0.90 (-0.06) 0.90 (-0.06) 0.90 (-0.06) 0.90 (-0.06) 0.96 (-0.00) 0.96 (-0.00) 0.96 (-0.00) 0.97 (+0.00) 0.97 (+0.00) 0.97 (+0.00) 0.97
Average 0.84 (-0.12) 0.90 (-0.06) 0.90 (-0.06) 0.91 (-0.04) 0.91 (-0.04) 0.91 (-0.04) 0.91 (-0.04) 0.94 (-0.01) 0.94 (-0.01) 0.94 (-0.01) 0.95 (-0.01) 0.95 (+0.00) 0.95 (+0.00) 0.95
DARPA APT Dataset [darpa]
FiveDirections 0.82 (+0.02) 0.80 (+0.00) 0.82 (+0.02) 0.82 (+0.02) 0.82 (+0.02) 0.82 (+0.02) 0.82 (+0.02) 0.82 (+0.02) 0.82 (+0.02) 0.82 (+0.02) 0.80 (+0.00) 0.82 (+0.02) 0.80 (+0.00) 0.80
Trace 0.87 (+0.00) 0.87 (+0.00) 0.87 (+0.00) 0.87 (+0.00) 0.93 (+0.07) 0.87 (+0.00) 0.93 (+0.07) 0.87 (+0.00) 0.87 (+0.00) 0.87 (+0.00) 0.93 (+0.07) 0.87 (+0.00) 0.87 (+0.00) 0.87
Theia 0.78 (-0.17) 0.78 (-0.17) 0.83 (-0.11) 0.83 (-0.11) 0.83 (-0.11) 0.83 (-0.11) 0.83 (-0.11) 0.83 (-0.11) 0.83 (-0.11) 0.83 (-0.11) 0.94 (+0.00) 0.94 (+0.00) 0.94 (+0.00) 0.94
Average 0.82 (-0.05) 0.82 (-0.06) 0.84 (-0.03) 0.84 (-0.03) 0.86 (-0.01) 0.84 (-0.03) 0.86 (-0.01) 0.86 (-0.01) 0.86 (-0.01) 0.86 (-0.01) 0.90 (+0.03) 0.90 (+0.03) 0.88 (+0.01) 0.89
Fileless Malware from [survivialism2021sp]
conhost 0.84 (-0.09) 0.89 (-0.04) 0.89 (-0.04) 0.89 (-0.04) 0.89 (-0.04) 0.89 (-0.04) 0.89 (-0.04) 0.89 (-0.04) 0.89 (-0.04) 0.89 (-0.04) 0.93 (+0.00) 0.93 (+0.00) 0.93 (+0.00) 0.93
explorer 0.72 (-0.26) 0.89 (-0.09) 0.91 (-0.07) 0.91 (-0.07) 0.89 (-0.09) 0.91 (-0.07) 0.89 (-0.09) 0.91 (-0.07) 0.91 (-0.07) 0.91 (-0.07) 0.98 (+0.00) 1.00 (+0.02) 1.00 (+0.02) 1.00
net 0.94 (-0.02) 0.94 (-0.02) 0.94 (-0.02) 0.94 (-0.02) 0.94 (-0.02) 0.94 (-0.02) 0.94 (-0.02) 0.91 (-0.06) 0.91 (-0.06) 0.91 (-0.06) 0.97 (+0.00) 0.97 (+0.00) 0.97 (+0.00) 0.97
reg 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 1.00 (+0.00) 1.00 (+0.00) 1.00 (+0.00) 1.00
wscript 0.89 (-0.07) 0.89 (-0.07) 0.89 (-0.07) 0.89 (-0.07) 0.89 (-0.07) 0.89 (-0.07) 0.89 (-0.07) 0.89 (-0.07) 0.89 (-0.07) 0.89 (-0.07) 0.95 (+0.00) 0.95 (+0.00) 0.95 (+0.00) 0.95
cscript 0.86 (-0.14) 0.86 (-0.14) 0.93 (-0.07) 0.93 (-0.07) 1.00 (+0.00) 0.93 (-0.07) 1.00 (+0.00) 1.00 (+0.00) 1.00 (+0.00) 1.00 (+0.00) 1.00 (+0.00) 1.00 (+0.00) 1.00 (+0.00) 1.00
mshta 0.85 (+0.00) 0.85 (+0.00) 0.85 (+0.00) 0.85 (+0.00) 0.85 (+0.00) 0.85 (+0.00) 0.85 (+0.00) 0.85 (+0.00) 0.85 (+0.00) 0.85 (+0.00) 1.00 (+0.15) 1.00 (+0.15) 0.85 (+0.00) 0.85
netsh 0.93 (-0.04) 0.91 (-0.07) 0.91 (-0.07) 0.89 (-0.09) 0.89 (-0.09) 0.89 (-0.09) 0.89 (-0.09) 0.89 (-0.09) 0.89 (-0.09) 0.89 (-0.09) 0.98 (+0.00) 0.98 (+0.00) 0.98 (+0.00) 0.98
python 0.58 (-0.41) 0.65 (-0.34) 0.65 (-0.34) 0.66 (-0.33) 0.64 (-0.35) 0.66 (-0.33) 0.64 (-0.35) 0.93 (-0.06) 0.93 (-0.06) 0.93 (-0.06) 0.99 (+0.00) 0.99 (+0.00) 0.99 (+0.00) 0.99
rundll32 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.96 (+0.00) 0.96 (+0.00) 0.96 (+0.00) 0.96
schtasks 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 1.00 (+0.00) 1.00 (+0.00) 1.00 (+0.00) 1.00
Average 0.86 (-0.10) 0.88 (-0.08) 0.89 (-0.08) 0.90 (-0.07) 0.90 (-0.07) 0.90 (-0.07) 0.90 (-0.07) 0.94 (-0.03) 0.94 (-0.03) 0.94 (-0.03) 0.97 (+0.00) 0.97 (+0.00) 0.97 (+0.00) 0.97
@@ A.3 Ablation Study (using GraphSAGE)
The ablation study utilizing the GraphSAGE  network revealed patterns consistent with those observed in the ablation study using GAT  network § 5.4. It was again noted that for Fileless Malware  datasets the graph structural features provided better feature selection than subgraph patterns. This pattern is observed because the malware graphs are distinctively different when compared against their benign counter part. Finally, the pattern that subgraph patterns with the graph structures provided the best feature selection.
A.4 ProvExplainer vs. SOTA Explainers
@@
Figure 7: Effectiveness of GraphSAGE  explainers in identifying prediction-relevant graph structures using $\text{fidelity}^{+}$  and $\text{fidelity}^{-}$
@@
Similar to the GAT -based results, ProvExplainer demonstrates superior performance across various datasets. The complexity of the provenance graphs, defined by their size and the nature of resource interactions, significantly influences the effectiveness of different explainers on specific datasets. For instance, on provenance graphs such as the APT graphs from [mukherjee2023sec], PGExplainer achieves the second-best performance. However, for the more complex or noisy DARPA APT graphs, GNNExplainer emerges as the second most effective method.
The structure of the Fileless Malware  dataset’s malware graphs differs notably from APT attacks, particularly due to the characteristics of attack campaigns. In this context, SubgraphX outperforms both GNNExplainer and PGExplainer. SubgraphX specializes in isolating distinctive malicious subgraphs that, when removed, significantly alter model predictions. Despite this strength, SubgraphX encounters limitations in scenarios where malware makes minimal, targeted changes to the system to evade detection, as observed in certain APT cases.
A.5 Case Studies: In-Depth Analysis
FiveDirections: Browser Extension. In the context of the described attack, where the drakon malware exploits the firefox.exe browser through a rogue extension (pass_mgr) the pattern #1 is created. Multiple instances of the malware process are created, each reading from the malicious files: pass_mgr.exe and passwordfile.dat. Additionally, they access essential dictionary files (en-US.aff) and cryptographic libraries (bcryptprimitives.dll) to maintain operational consistency, allowing them to conduct their malicious activities efficiently. The pattern #2 captures the data extraction behavior through sensitive file accesses. These malware access sensitive information and system configuration files like WindowsShell.Manifest, and wintrust.dll, along with initial malware files (addons, tzres.dll, userenv.dll).
@@ A.5 Case Studies: In-Depth Analysis
FiveDirections: Copykatz. The pattern #2 starts its kill chain stage of initial access [apts1] by writing and executes a malware masquerading as firefox.exe which contains the Mimikatz and Copykatz modules. ntdll.dll [ntdll] includes multiple kernel-mode functions which enables the “Windows Application Programming Interface (API)” and bcryptprimitives.dll [bcryptprimitives] contain functions implementing cryptographic primitives, which are essential for Mimikatz [mimikatz] and Copykatz.
The firefox.exe malware write its payload into temporary files, such as virtuous and tropical. These files are then executed by malware in the system creating pattern #5. Subsequently, another malicious instance of the firefox.exe malware initiates a chain or read-execute cycle by executing the malware template as well as making external IP connections. This dependency correlation between the malware and its parent is characterized by inheritance, where the children require the same system library files as the parent to function correctly. The DLLs involved are read from ProgramFiles and System32.
GNNExplainer and PGExplainer correctly identified the data extraction stage where four malicious firefox.exe processes make C2 connections to the external IPs (202.179.137.58 and 217.160.205.44). firefox.exe was invoked by malware masquerading as runtimebroker.exe, so if GNNExplainer masks out the process creation edge of the children firefox.exe, that would lead to a graph cut with two separate graphs, leading to a change in the prediction. SubgraphX also highlighted similar firefox.exe processes that are created by the first runtimebroker.exe, but instead of making external C2 connection it created another firefox.exe process for rendering content from localhost (127.0.0.1).
Trace: Phishing E-mail. The pattern #2 captured the staging behavior of the Trojan. The Trojan was downloaded from www.nasa.ng, executed and replicated itself within the system. Specifically, the malware named nasa.ng is placed in /home/admin/.mozilla/firefox/ and /usr/local/firefox-54.0.1/obj-x86_64-pc-linux-gnu/. The Trojan created new malware with benign names such as firefox to effectively evade detectors, to replicate unhindered and overloaded the system with malicious processes. After the malware successfully staged, the malware masquerading as /bin/sh to read the malicious script staged in (/etc/update-motd.d/00-header/) and executed it to create multiple copies of itself. The malware reads various system configuration files present in /etc/protocols/, /etc/lsb-release/, and /etc/hosts.deny/. Reading sensitive system configuration files are essential to build the system profile. The pattern #5 captured the dependency correlation of the malware created that inherited configurations.
@@
GNNExplainer was able to correctly capture the staging behavior where the malware from nasa.ng read shared library (/usr/lib/x86_64-linux-gnu/*.so.*) and cache file (/usr/share/applications/mimeinfo.cache, /usr/lib/x86_64-linux-gnu/*/loaders.cache). PGExplainer incorrectly indicated benign substructures, but SubgraphX correctly captured the inheritance behavior of firefox.exe executing multiple times with the argument file http://www.nasa.ng/, to create the malware clones from the template. SOTA explainers missed the malware’s ultimate goal of reading sensitive files, which was only captured by ProvExplainer.

### v7

@@ IX Conclusion
We introduced ProvExplainer, a framework designed to address a critical gap in explainable ML for cybersecurity. It provides transparent, interpretable insights into GNN-based IDS in the system provenance domain. By leveraging security-aware discriminative subgraph patterns and graph structural features, ProvExplainer facilitates instance-level explanations that consider both local and global graph properties. This dual-level interpretability is essential for removing the opacity and help with the adoption of GNN-based IDS. ProvExplainer is a critical advancement in explainable AI for the security domain and a scalable framework for real-world deployment.
@@ Appendix A Appendix
Table III: Node and edge types for in-house system provenance with associated attributes.
Types Attributes
Nodes
process signature, executable name, pid
file owner (uid, gid), name, inode
socket (IP) dstip, srcip, dstport, srcport, type
Edges
process $\rightarrow$ process command args, starttime
process $\rightarrow$ file read, write, amount
process $\rightarrow$ IP address send, recv, amount
A-A Implementation
Our data collection module can operate on Windows systems by using the Windows ETW [72] and on Linux system by using Linux auditd [73] frameworks to collect relevant system calls regarding files, processes and network sockets. These include system calls for (1) file operations (e.g., read(), write(), unlink()), (3) network socket operations (e.g., connect(), accept()), (4) process operations (e.g., create(), exec(), and exit()). The system-level data is stored in a PostgresSQL database. Table III summarizes the event collection schema for our in-house deployment. While it includes major nodes and event types annotated with textual and numerical attributes, detailed entries of the event collection differ across OS environment.
ProvExplainer is implemented using python and graph generation framework is implemented in java. The GNN training and evaluation pipelines are implemented using Deep Graph Library (DGL) [58] framework, and the surrogate DTs are implemented using sklearn [74]. We used the frequent subgraph miner, GERM [39], to mine patterns with a minimum support of 1 and a maximum subgraph size of 10. The discriminant subgraph patterns were selected by maximizing the scoring function. The scoring function’s hyperparameters and DT were empirically tuned using grid search by iteratively evaluating ProvExplainer’s performance using Fidelity+/- and Precision/Recall metric on cross-validation folds.
A-B Dataset Statistics
@@
Applications
# of Benign
Graphs
# of Anomaly
Graphs
Avg # of Benign
Nodes / Edges
Avg # of Anomaly
Nodes / Edges
@@
Fileless Malware from [47]
@@
Anomaly Detection Dataset. The dataset statistics for the anomaly detection dataset is seen in Table IV. The Fileless Malware samples were downloaded from [30] and selected from recent studies, [75, 70, 71, 47]. The benign provenance graphs for the anomaly detection dataset were sourced from the DARPA dataset, which contained 5,695 benign graphs with an average of 669.81 vertices and 982.18 edges (Table IV); the Fileless Malware dataset, which contained 19,422 benign graphs with an average of 60.44 vertices and 54.15 edges; and the APT dataset, which contained 6,291 benign graphs with an average of 55.94 vertices and 83.29 edges. The corresponding anomaly graphs are sourced from the same datasets: The DARPA dataset contained 30 anomalous graphs with an average of 777.08 vertices and 1011.19 edges (Table IV); the Fileless Malware contained 1,043 anomalous graphs with an average of 63.72 vertices and 89.93 edges; and the APT dataset contained 2,928 graphs with an average of 58.38 vertices and 51.69 edges.
A-C Ablation Study (using GraphSAGE)
(a) DARPA APT Dataset [darpa]
(b) APT Dataset from [46]
(c) Fileless Malware from [47]
Figure 7: Agreement (higher is better) measured using WMA F1 of surrogate decision trees versus the GraphSAGE model across different feature sets.
@@ A-C Ablation Study (using GraphSAGE)
The ablation study utilizing the GraphSAGE network revealed patterns consistent (as shown in Figure 7) with those observed in the ablation study using GAT network § VI-E. It was again noted that for Fileless Malware datasets the graph structural features provided better feature selection than subgraph patterns. This pattern is observed because the malware graphs are distinctively different when compared against their benign counter part. Finally, the pattern that subgraph patterns with the graph structures provided the best feature selection.
A-D ProvExplainer vs. SOTA Explainers
@@
Figure 8: Effectiveness of GraphSAGE explainers in identifying prediction-relevant graph structures using $\text{fidelity}^{+}$ and $\text{fidelity}^{-}$ .
@@
Similar to the GAT -based results, ProvExplainer demonstrates superior performance across various datasets. The complexity of the provenance graphs, defined by their size and the nature of resource interactions, significantly influences the effectiveness of different explainers on specific datasets. For instance, on provenance graphs such as the APT graphs from [46], PGExplainer achieves the second-best performance. However, for the more complex or noisy DARPA APT graphs, GNNExplainer emerges as the second most effective method.
A-E Case Study: FiveDirections (Copykatz) [darpa]
(a) Attack Graph.
(b) SOTA GNN explainations.
(c) Pattern #2.
Description. In a sophisticated attack, a hijacked version of usdoj.gov exploits Firefox to deploy drakon malware to the victim host. Then, drakon uses the elevate driver to escalate privileges and masquerade as the runtimebroker.exe system program. Finally, the malicious runtimebroker.exe instance connects to a command and control (C2) server to download and execute Copykatz (an older version of Mimikatz) to harvest and exfiltrate host credentials, as illustrated in Figure 9.
Features: Degree centrality and pattern #2.
System Interpretation. The pattern shown in 9(c) highlights a multi-stage attack sequence characteristic of the Firefox Drakon APT focusing on elevating the privilege using Copykatz. Process P3 (firefox.exe) initiates the attack by exploiting its backdoor to download the payload, File F1 (ctfhost2.exe) that contains the drakon malware. This payload is subsequently executed by Process P1 (runtimebroker.exe), which modifies critical system libraries, File F2 (ntdll.dll) and File F3 (kernel32.dll). These modifications indicate the preparation of the system environment for privilege escalation and sustained malicious operations. Process P2 (svchost.exe) interaction with File F2 emphasizes its role in leveraging system-level components to propagate the attack, consistent with the exploitation of driver signing for privilege elevation.
The graph structure captures the interdependencies between processes and system-critical files, illustrating the deliberate flow of the attack. By chaining interactions, such as runtimebroker.exe modifying ntdll.dll and svchost.exe subsequently accessing it, the graph provides insights into how the malware manipulates system operations. Furthermore, the reuse of Firefox as both an initial access vector and a conduit for kernel32.dll interactions highlights the advanced tactics employed to achieve persistence and exploitation. These observations align closely with the described TTPs of the Firefox Drakon Copykatz APT, offering a clear mapping of graph elements to specific stages of the attack lifecycle. The attack graph contains high quantities of file read and writes flowers for a relatively small number of process nodes acting as bridges between the flowers, so it leads to a high degree of centrality.
ProvExplainer vs. SOTA explainers. GNNExplainer and PGExplainer pinpoint the stage where multiple Firefox processes connect to the C2 servers. This activity traces back to the malicious runtimebroker instance. Notably, SubgraphX detects an alternate trajectory where a Firefox process, instead of reaching out to external servers, spawns another process aimed at local content manipulation, showcasing the malware’s versatility in engaging with both external and internal resources for its objectives. While SOTA explainers have demonstrated proficiency in identifying the final stages of this data breach, only ProvExplainer effectively captured both the initial infection and its privilege escalation attempts. This distinction highlights the importance of recognizing early-stage indicators for root cause analysis.
A-F In-Depth Analysis of Case Studies
Trace: Phishing E-mail [darpa]. The pattern #2 captured the staging behavior of the Trojan. The Trojan was downloaded from www.nasa.ng, executed and replicated itself within the system. Specifically, the malware named nasa.ng is placed in /home/admin/.mozilla/firefox/ and /usr/local/firefox-54.0.1/obj-x86_64-pc-linux-gnu/. The Trojan created new malware with benign names such as firefox to effectively evade detectors, to replicate unhindered and overloaded the system with malicious processes. After the malware successfully staged, the malware masquerading as /bin/sh to read the malicious script staged in (/etc/update-motd.d/00-header/) and executed it to create multiple copies of itself. The malware reads various system configuration files present in /etc/protocols/, /etc/lsb-release/, and /etc/hosts.deny/. Reading sensitive system configuration files are essential to build the system profile. The pattern #5 captured the dependency correlation of the malware created that inherited configurations.
@@ A-F In-Depth Analysis of Case Studies
GNNExplainer was able to correctly capture the staging behavior where the malware from nasa.ng read shared library (/usr/lib/x86_64-*/*.so.*) and cache file (/usr/share/applications/mimeinfo.cache, /usr/lib/x86_64-*/*/loaders.cache). PGExplainer incorrectly indicated benign substructures, but SubgraphX correctly captured the inheritance behavior of firefox.exe executing multiple times with the argument file http://www.nasa.ng/, to create the malware clones from the template. SOTA explainers missed the malware’s ultimate goal of reading sensitive files, which was only captured by ProvExplainer.
FiveDirections: Browser Extension [darpa]. In the context of the described attack, where the drakon malware exploits the firefox.exe browser through a rogue extension (pass_mgr) the pattern #1 is created. Multiple instances of the malware process are created, each reading from the malicious files: passwordfile.dat and pass_mgr.exe. Additionally, they access essential dictionary files (en-US.aff) and cryptographic libraries (bcryptprimitives.dll ) to maintain operational consistency, allowing them to conduct their malicious activities efficiently. The pattern #2 captures the data extraction behavior through sensitive file accesses. These malware access sensitive information and system configuration files like WindowsShell.Manifest, and wintrust.dll, along with initial malware files (addons, tzres.dll, userenv.dll).
@@
FiveDirections: Copykatz [darpa]. The pattern #2 starts its kill chain stage of initial access [76] by writing and executes a malware masquerading as firefox.exe which contains the Mimikatz and Copykatz modules. ntdll.dll [77] includes multiple kernel-mode functions which enables the “Windows Application Programming Interface (API)” and bcryptprimitives.dll [78] contain functions implementing cryptographic primitives, which are essential for Mimikatz [40] and Copykatz.
The firefox.exe malware write its payload into temporary files, such as ctfhost.exe, virtuous and tropical. These files are then executed by malware in the system creating pattern #5. Subsequently, another malicious instance of the firefox.exe malware initiates a chain or read-execute cycle by executing the malware template as well as making external IP connections. This dependency correlation between the malware and its parent is characterized by inheritance, where the children require the same system library files as the parent to function correctly. The DLLs involved are read from ProgramFiles and System32.
GNNExplainer and PGExplainer correctly identified the data exfiltration stage where four malicious firefox.exe make C2 connections to the external IPs (202.179.137.58 and 217.160.205.44). firefox.exe was invoked by malware masquerading as a popular system program, runtimebroker.exe. GNNExplainer correctly masks out the process creation edges of the children firefox.exe, to create disjoint graphs and leading to prediction change. SubgraphX also highlighted firefox.exe processes that are created by runtimebroker.exe, but instead of making external connection it creates other firefox.exe for rendering content from localhost (127.0.0.1).